# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 4.4 Action path stochastic decision forests
    and $g\colon D\to\{0,1\}$ a map. If $D\neq\emptyset$, then we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies assumptions AP.C0, AP.C1, AP.C2. Moreover, assumption AP.C3 is satisfied.
@@
It is important to note that a choice $c=c(A_{<t},i,g)\in\mathscr{C}_{t}$ as above, available at a random move ${\bm{\mathsf{x}}}$, with $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable $g\colon D\to\mathbb{A}^{i}$, $D\in\mathscr{A}$, can condition on two distinct things: first, on the endogenous information $A_{<t}$, and second, on the exogenous information via the $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurability of $g$. For instance, one can the imagine the case in that there are ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}\setminus\{{\bm{\mathsf{x}}}\}$ with $\mathscr{F}^{i}_{\bm{\mathsf{x}}}=\mathscr{F}^{i}_{{\bm{\mathsf{x}}}^{\prime}}$ and $\mathfrak{t}({\bm{\mathsf{x}}})=t=\mathfrak{t}({\bm{\mathsf{x}}}^{\prime})$, $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable $g^{\prime}\colon D\to\mathbb{A}^{i}$ with $g\neq g^{\prime}$ and $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ such that $c(A^{\prime}_{<t},i,g^{\prime})$ is available at ${\bm{\mathsf{x}}}^{\prime}$ and an agent’s complete contingent plan of action specifies $c(A_{<t},i,g)$ at ${\bm{\mathsf{x}}}$, but $c(A^{\prime}_{<t},i,g^{\prime})$ at ${\bm{\mathsf{x}}}^{\prime}$. In the context of example 3.28, with ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f)$ and ${\bm{\mathsf{x}}}^{\prime}={\bm{\mathsf{x}}}_{t}(f^{\prime})$ for suitable $f,f^{\prime}\in\mathbb{A}^{\mathbb{T}}$, this would mean that, even if $Y_{{\bm{\mathsf{x}}}_{u}(f)}$, $u\leq t$, and $Y_{{\bm{\mathsf{x}}}_{u}(f^{\prime})}$, $u\leq t$, generate the same $\sigma$-algebra, i.e., $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f^{\prime})}$, an agent need not make the very choice the agent makes at ${\bm{\mathsf{x}}}_{t}(f)$ at the counterfactual random move ${\bm{\mathsf{x}}}_{t}(f^{\prime})$ as well.
@@
Let us close on a note about this specific example for that we consider an agent only having choices of the form $c=c(A_{<t},i,g)$ as in the preceding paragraph. In case $A_{<t}=\mathbb{A}^{[0,t)_{\mathbb{T}}}$ for all choices available to this agent, and moreover $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{G}_{t}$ for all $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in D_{t,f}$, and a fixed filtration $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ (case 1 in subsection 3.4), then the complete contingent plans of action this agent can build, model what is described by the term “open loop” in the control- and game-theoretic literature. If, conversely, only choices with singleton $A_{<t}$ are available to the agent and $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}$ depends on a stochastic state controlled by $f$, for $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in D_{t,f}$, in the sense of case 2 of subsection 3.4, possibly in combination with example 3.28, then we obtain what is subsumed under the term “closed loop” (see, e.g., [20, 21]).
@@
Note that there are mixed regimes. “Large” alias uninformative $A_{<t}$ and $\mathscr{F}^{i}$ like in case 2 of subsection 3.4 may coexist as well, which is often also called “closed loop” in the control and differential games literature (see, e.g., the corresponding discussion in [13, p. 72–76]), but actually the complete contingent plans of action the agent can make only close the loop with respect to exogenous information: the agent can only react to the controlled noise, but not explicitly to the employed controls themselves which can make a crucial difference when evaluating counterfactuals, especially in case several agents interact. On the other hand, “small” alias informative $A_{<t}$ may coexist with case 1 of subsection 3.4, that is, “$\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{G}_{t}$”, which is near, or in the extreme case of singleton $A_{<t}$ equal to closed loops with respect to endogenous information, but is equivalent, on the exogenous information side, to the ignorance of any controlled stochastic state. In any case, we conclude, that the terminology of “closed” and “open” loops must be used with respect to both endogenous and exogenous information. It requires some caution in the mixed regimes just discussed, especially if the underlying stochastic decision forest, exogenous information structures and adapted choices are not properly specified.
@@ Conclusion
It is possible to implement general stochastic processes as background noise on refined partitions-based decision trees without running into outcome generation problems for a “nature” agent and allowing for a rigorous decision-theoretic interpretation of the relation of endogenous and exogenous information and choices. This is an improvement over the existing state of the art in both refined partitions-based decision and game theory in [5, 1], and stochastic control and differential games theory in [34, 29, 13, 14]. This has been achieved by abandoning the assumption of a “nature” agent, and instead constructing a theory of stochastic decision forests. These satisfy duality between outcomes and nodes and can be represented as forests of decision trees. They allow for a notion of similarity across moves on different trees, given by random moves, which necessarily form the moves of a decision tree in its own right. They serve as the basis for both a structure of exogenous information revelation similar to filtrations in probability theory, and a concept of adapted choices compatible with exogenous information structures which capture measurability assumptions on choices typically made in the literature and also shed light on the decision-theoretic meaning of open and closed loop controls in the stochastic setting.
@@
In addition, a general model of action path stochastic decision forests has been constructed, given in terms of a small number of easily verifiable conditions. This model can serve as a unified decision-theoretic basis for a large class of stochastic decision problems, for instance in discrete time. Moreover, as will be shown in the third paper, it constitutes one step towards explaining stochastic decision problems in continuous time approximately.
@@
In order to obtain a theory of stochastic extensive forms, encompassing the classical theory in [5], it still remains to formulate consistency criteria on stochastic decision forests equipped with a set of agents, each agent provided with an exogenous information structure, a reference choice structure and a set of adapted choices, to define strategies in the classical decision-theoretic sense going back to Savage, as prepared in this paper, and to ask the question whether and, if yes, how outcome existence and uniqueness (well-posedness) for strategy profiles can be characterised for stochastic extensive forms. Moreover, the question, implicit in the present paper, can then be formally addressed which stochastic extensive forms can be faithfully represented as classical extensive forms, that is, over a rooted decision tree, with a “nature” agent simulating the noise. An answer to this can formally substantiate the claim that stochastic extensive forms, and sdfs which they are based on, form a true and useful generalisation of the existing theory. Furthermore, we note in a spirit related to [35] that in order to allow for subgame-perfect equilibrium analysis, a decision-theoretically explainable and general model of subgames in stochastic extensive form decision problems needs to be established. All this constitutes the programme for the second paper of the present series.
@@ Acknowledgements
First and foremost, I am particularly indebted to my doctoral supervisor Christoph Knochenhauer who read and commented on several earlier versions of this paper, encouraged me to pursue this project and actively supported me in presenting its contents at workshops. Special thanks are due to Frank Riedel for the helpful discussion we had in Berlin in 2022, especially for giving useful hints regarding the game-theoretic literature. I also want to thank Elizabeth Baldwin, Miguel Ballester, and Samuel N. Cohen for inspiring discussions on decision and control theory during my stay in Oxford in 2023. For their questions and comments, I thank the attendants of the workshops and seminars I had the opportunity to present and discuss earlier versions of this particular project at, in Oxford, Bielefeld, Kiel, Berlin, Palaiseau, including Daniel Andrei, Karolina Bassa, Peter E. Caines, Fanny Cartellier, Sebastian Ertel, Ali Lazrak, Kristoffer Lindensjö, Christopher Lorenz, Berenice Anne Neumann, Manos Perdikakis, Jan-Henrik Steg, Peter Tankov, and Jacco Thijssen. This project was partially funded by Deutsche Forschungsgemeinschaft (DFG) – Project ID 410208580 – IRTG2544 (“Stochastic Analysis in Interaction”).
@@ A.1 Section 1
We now prove the statement of the second sentence. Suppose the forest $(F,\geq)$ to be rooted, and let $T\in\mathcal{F}$. By definition of $\mathcal{F}$, $T$ is non-empty. Let $x\in T$. Then the principal up-set $\uparrow x=\{y\in F\mid y\geq x\}$ in $F$ contains a maximal element $y$. By definition of $T$, $y\in\uparrow x$ implies $y\in T$. Futhermore, let $z\in T$ with $z\geq y$. As $z\in F$, the maximality of $y$ implies $z=y$. Hence, $y$ is maximal in $(T,\geq)$ as well. Therefore, $(T,\geq)$ is rooted according to our definition.
It remains to prove the existence of a maximum of $(T,\geq)$. $T$ is non-empty, hence there is $x\in T$. By assumption, there is a maximal element $y\in\uparrow x$. $\uparrow x$ is a chain, because $(F,\geq)$ is a forest. Hence, $y$ is a maximum of $\uparrow x$ with respect to the induced order. Let $z\in T$ be an arbitrary element of $T$. Then $\uparrow x\cap\uparrow z\neq\emptyset$. Take an element $u$ in this intersection, which is also an element of $T$, because $u\sim x$. Then $y\geq u\geq z$, whence $y\geq z$. Thus $y$ is a maximum of $(T,\geq)$. ∎
@@
The second part follows from the first, using that all maximal chains $w$ in $(T,\supseteq)$ satisfy $V_{T}\in w$. Indeed, as a maximal chain $w$ is non-empty, and any element $x\in w\subseteq T$ satisfies $V_{T}\supseteq x$ by definition of $V_{T}$. By maximality of the chain $w$, we infer $V_{T}\in w$. ∎
@@
Let $F$ be a decision forest over a set $V$ with $f$ as in definition 1.3. Let $\mathcal{F}$ be the set of $(F,\supseteq)$’s connected components. Then, for any $v\in V$ and any $T\in\mathcal{F}$, we have $f(v)\subseteq T$ iff $v\in V_{T}$.
@@
Let $v\in V$ and $T\in\mathcal{F}$. By proposition 1.4, $v\in V_{T}$ iff $V_{T}\in f(v)$. By lemma A.42, part (2), this is equivalent to $f(v)\subseteq T$. ∎
@@
(Ad necessity):  Suppose $F$ to be a decision forest over $V$. Let $W$ be the set of maximal chains in $(F,\supseteq)$, and $f\colon V\rightarrow W$ a bijection such that, for every $x\in F$, $({\mathcal{P}}f)(x)=W(x)$, just as in definition 1.3.
We claim that for all $v\in V$ there is a unique $T\in\mathcal{F}$ such that $f(v)\subseteq T$. The claim (1) then follows with lemma A.43.
For the proof, let $v\in V$. As the maximal chain $f(v)$ is non-empty, there is $x\in f(v)$. As $\mathcal{F}$ is a partition of $F$, there is a unique $T\in\mathcal{F}$ with $x\in T$. For arbitrary $y\in f(v)$, we have $x\subseteq y$ or $y\subseteq x$, since $f(v)$ is a chain. Hence, $x\sim y$, or in other words $y\in T$. Thus $T$ is the unique element of $\mathcal{F}$ with $f(v)\subseteq T$.
For claim (2), let $T\in\mathcal{F}$. Then $(T,\supseteq)$ defines a rooted tree with root $V_{T}$, by definition of $\mathcal{F}$. Moreover, for all $v\in V$, we have $f(v)\in W(V_{T})$ iff $v\in V_{T}$, by proposition 1.4. Hence, $f_{T}:=f|_{V_{T}}$ yields a bijection $V_{T}\rightarrow W(V_{T})$.
By lemma A.42, part (3), second sentence, $W(V_{T})$ equals the set of maximal chains in $(T,\supseteq)$. By definition of $V_{T}$, all $x\in T$ satisfy $V_{T}\supseteq x$ so that the of definition of $f$ yields:
@@
By lemma A.42, part (3), first sentence, $W(x)$ consists exactly of all maximal chains in $(T,\supseteq)$ containing $x$. Hence, $T$ is its own representation by decision paths.
(Ad sufficiency):  Let $F$ satisfy properties (1) and (2) from the theorem. It remains to show that $F$ is its own representation by decision paths, i.e., part (2) in definition 1.3.
By assumption (2) and part (3) of lemma A.42, for each $T\in\mathcal{F}$, there is a bijection $f_{T}\colon V_{T}\rightarrow W(V_{T})$ such that for all $x\in T$, $(\mathcal{P}f_{T})(x)$ equals the set of decision paths in $(T,\supseteq)$ containing $x$. By part (3) of lemma A.42, this set is equal to $W(x)$. By assumption (1), we obtain a map $f\colon V\rightarrow W$ satisfying, for any $T\in\mathcal{F}$ and $v\in V_{T}$, $f(v)=f_{T}(v)$. $f$ is injective by lemma A.42, part (2), since $\mathcal{F}$ is a partition. $f$ is also surjective, hence bijective, since every maximal chain $w\in W$ contains an element $x$ for that there is $T\in\mathcal{F}$ with $x\in T$. Hence $w\in W(x)$ which is a subset of $W(V_{T})$ by lemma A.42, part (3). Hence, there is $v\in V_{T}$ with $f(v)=f_{T}(v)=w$.
Let $x\in F$. Then there is $T\in\mathcal{F}$ with $x\in T$. Then $V_{T}\supseteq x$. Hence, $({\mathcal{P}}f)(x)=({\mathcal{P}}f_{T})(x)$. By definition of $f_{T}$ and part (3) of lemma A.42, we have
@@
Lemma A.42, part (3), implies that the latter set equals $W(x)$. Thus, $(\mathcal{P}f)(x)=W(x)$. ∎

### v2

@@ 4.4 Action path stochastic decision forests
    we have $D\neq\emptyset$. Further, let $g\colon D\to\{0,1\}$ be a map and let $A_{<t}=\{1\}^{[0,t)}$. Then, we have $c(A_{<t},i,g)\in\mathscr{C}_{t}$, that is, it satisfies Assumptions AP.C0, AP.C1, AP.C2. Moreover, if for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$, Assumption AP.C3 is satisfied for $(A_{<t},i,g)$ and $\mathscr{C}^{i}$.
The proofs of these claims can be found in the corresponding part of the appendix.
@@
It is important to note that a choice $c=c(A_{<t},i,g)\in\mathscr{C}_{t}$ as above, available at a random move ${\bm{\mathsf{x}}}$, with $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable $g\colon D\to\mathbb{A}^{i}$, $D\in\mathscr{E}$, can condition on two distinct things: first, on the endogenous information $A_{<t}$, and second, on the exogenous information via the $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurability of $g$. For instance, one can the imagine the case in that there are ${\bm{\mathsf{x}}}^{\prime}\in\tilde{\bm{\mathsf{X}}}^{i}\setminus\{{\bm{\mathsf{x}}}\}$ with $\mathscr{F}^{i}_{\bm{\mathsf{x}}}=\mathscr{F}^{i}_{{\bm{\mathsf{x}}}^{\prime}}$ and $\mathfrak{t}({\bm{\mathsf{x}}})=t=\mathfrak{t}({\bm{\mathsf{x}}}^{\prime})$, $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}}$-measurable $g^{\prime}\colon D\to\mathbb{A}^{i}$ with $g\neq g^{\prime}$ and $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ such that $c(A^{\prime}_{<t},i,g^{\prime})$ is available at ${\bm{\mathsf{x}}}^{\prime}$ and an agent’s complete contingent plan of action specifies $c(A_{<t},i,g)$ at ${\bm{\mathsf{x}}}$, but $c(A^{\prime}_{<t},i,g^{\prime})$ at ${\bm{\mathsf{x}}}^{\prime}$. In the context of Example 3.31, with ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f)$ and ${\bm{\mathsf{x}}}^{\prime}={\bm{\mathsf{x}}}_{t}(f^{\prime})$ for suitable $f,f^{\prime}\in\mathbb{A}^{\mathbb{T}}$, this would mean that, even if $Y_{{\bm{\mathsf{x}}}_{u}(f)}$, $u\leq t$, and $Y_{{\bm{\mathsf{x}}}_{u}(f^{\prime})}$, $u\leq t$, generate the same $\sigma$-algebra, i.e. $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f^{\prime})}$, an agent need not make the very choice the agent makes at ${\bm{\mathsf{x}}}_{t}(f)$ at the counterfactual random move ${\bm{\mathsf{x}}}_{t}(f^{\prime})$ as well.
@@
Let us close on a note about this specific example for that we consider an agent only having choices of the form $c=c(A_{<t},i,g)$ as in the preceding paragraph. In case $A_{<t}=\mathbb{A}^{[0,t)_{\mathbb{T}}}$ for all choices available to this agent, and moreover $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{G}_{t}$ for all $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in D_{t,f}$, and a fixed filtration $\mathscr{G}=(\mathscr{G}_{t})_{t\in\mathbb{T}}$ (case 1 in Subsection 3.4), then the complete contingent plans of action this agent can build, model what is described by the term “open loop” in the control- and game-theoretic literature. If, conversely, only choices with singleton $A_{<t}$ are available to the agent and $\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}$ depends on a stochastic state controlled by $f$, for $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in D_{t,f}$, in the sense of case 2 of Subsection 3.4, possibly in combination with Example 3.31, then we obtain what is subsumed under the term “closed loop” (see, e.g. [20, 21]).
@@
Note that there are mixed regimes. “Large” alias uninformative $A_{<t}$ and $\mathscr{F}^{i}$ like in case 2 of Subsection 3.4 may coexist as well, which is often also called “closed loop” in the control and differential games literature (see, e.g. the corresponding discussion in [13, p. 72–76]), but actually the complete contingent plans of action the agent can make only close the loop with respect to exogenous information: the agent can only react to the controlled noise, but not explicitly to the employed controls themselves which can make a crucial difference when evaluating counterfactuals, especially in case several agents interact. On the other hand, “small” alias informative $A_{<t}$ may coexist with case 1 of Subsection 3.4, that is, “$\mathscr{F}^{i}_{{\bm{\mathsf{x}}}_{t}(f)}=\mathscr{G}_{t}$”, which is near, or in the extreme case of singleton $A_{<t}$ equal to closed loops with respect to endogenous information, but is equivalent, on the exogenous information side, to the ignorance of any controlled stochastic state. In any case, we conclude, that the terminology of “closed” and “open” loops must be used with respect to both endogenous and exogenous information. It requires some caution in the mixed regimes just discussed, especially if the underlying stochastic decision forest, exogenous information structures and adapted choices are not properly specified.
@@ Conclusion
It is possible to implement general stochastic processes as background noise on refined partitions-based decision forests without encountering outcome generation problems for a “nature” agent, while allowing for a rigorous decision-theoretic interpretation of the relationship between endogenous and exogenous information and choices. This represents an improvement over the existing state of the art in both refined partitions-based decision and game theory (e.g. in [5, 1]) and stochastic control and differential games theory (e.g. [38, 33, 13, 14]). This has been achieved by abandoning the assumption of a “nature” agent and instead constructing a theory of stochastic decision forests. These decision forests satisfy duality between outcomes and nodes and can be represented as forests of decision trees. They allow for a notion of similarity across moves on different trees, given by random moves, which, under weak hypotheses on the underlying order-theoretic structure, form the moves of a decision tree in its own right. They serve as the basis for both a structure of exogenous information revelation similar to filtrations in probability theory and a concept of adapted choices compatible with exogenous information structures. Adapted choices capture measurability assumptions on choices typically made in the literature and also provide insight into the decision-theoretic meaning of open and closed loop controls in the stochastic setting.
@@
In addition, a general model of action path stochastic decision forests has been constructed, based on a small number of easily verifiable conditions. This model can serve as a unified decision-theoretic foundation for a large class of stochastic decision problems, including those in discrete time. Moreover, as will be shown in the third paper, it constitutes a step toward approximately explaining the extensive form characteristics of stochastic decision problems in continuous time.
@@
In order to develop a theory of stochastic extensive forms that encompasses the classical theory in [5], it is still necessary to formulate consistency criteria on stochastic decision forests equipped with a set of agents, each agent provided with an exogenous information structure, a reference choice structure, and a set of adapted choices. This would allow us to define strategies as Savage acts, as prepared in this paper, and to address the question of whether and, if so, how outcome existence and uniqueness for strategy profiles – and in that sense, essentially, “well-posedness” – can be characterised for stochastic extensive forms. Based on this, we can address the question, implicit in the present paper, of which stochastic extensive forms can and which cannot be faithfully represented as extensive forms with an additional “nature” agent that, by executing a strategy, generates exogenous information as endogenous information. Furthermore, we note, in a spirit related to [41], that in order to allow for subgame-perfect or perfect Bayesian equilibrium analysis, a decision-theoretically explainable and general model of subgames, histories, and information sets in stochastic extensive form decision problems needs to be established. Moreover, such an analysis requires an abstract concept of preferences, e.g. via beliefs and utility functions, following entirely from basic decision-theoretic principles. Last but not least, such a theory is expected to enable the construction and study of stochastic extensive forms based on action path data, providing a large class of game-theoretic models and, especially, the basis of an approximation theory for stochastic differential games. This remains to be shown as well. Together, all of this constitutes the programme for the second paper [40] of the present series.
@@ Acknowledgements
First and foremost, the author is particularly indebted to his doctoral supervisor, Christoph Knochenhauer, who read and commented on several earlier versions of this paper, encouraged him to pursue this project, and actively supported him in presenting its contents at workshops. Special thanks are due to Frank Riedel for the helpful discussion we had in Berlin in 2022, especially for providing useful insights regarding the game-theoretic literature. The author also wants to thank Elizabeth Baldwin, Miguel Ballester, and Samuel N. Cohen for inspiring discussions on decision and control theory during his stay in Oxford in 2023. The author is also grateful to the attendants of the workshops and seminars where he had the opportunity to present and discuss earlier versions of this particular project, in Oxford, Kiel, Berlin, Palaiseau, including Daniel Andrei, Karolina Bassa, Peter E. Caines, Fanny Cartellier, Sebastian Ertel, Ali Lazrak, Kristoffer Lindensjö, Christopher Lorenz, Berenice Anne Neumann, Manos Perdikakis, Jan-Henrik Steg, Peter Tankov, and Jacco Thijssen, for their questions and comments. Partial funding by Deutsche Forschungsgemeinschaft (DFG) through, first, the IRTG 2544 Stochastic Analysis in Interaction, Project ID: 410208580, and, second, under Germany’s Excellence Strategy, the EXC 2046/1 The Berlin Mathematics Research Center MATH+, Project ID: 390685689, is gratefully acknowledged.
@@ A.1 Section 1
We now prove the statement of the second sentence. Suppose the forest $(F,\geq)$ to be rooted, and let $T\in\mathcal{F}$. By definition of $\mathcal{F}$, $T$ is a non-empty tree. Let $x\in T$. Then the principal up-set $\uparrow x=\{y\in F\mid y\geq x\}$ in $F$ contains a maximal element $y$. By definition of $T$, $y\in\uparrow x$ implies $y\in T$. $\uparrow x$ is a chain, because $(F,\geq)$ is a forest. Hence, $y$ is even a maximum of $\uparrow x$ with respect to the induced order. Let $z\in T$ be an arbitrary element of the tree $T$. Then $T\cap\uparrow x\cap\uparrow z\neq\emptyset$. Take an element $u$ in this intersection. Then $y\geq u\geq z$, whence $y\geq z$. Thus $y$ is a maximum of $(T,\geq)$. In particular, $y$ is maximal in $(T,\geq)$. Therefore, $(T,\geq)$ is a rooted tree according to our definition and has a maximum. ∎
@@
The second part follows from the first, since all maximal chains $w$ in $(T,\supseteq)$ satisfy $V_{T}\in w$. Indeed, as a maximal chain in the rooted tree $(T,\supseteq)$, $w$ is non-empty, and any element $x\in w\subseteq T$ satisfies $V_{T}\supseteq x$ by definition of $V_{T}$. By maximality of the chain $w$, we infer $V_{T}\in w$. ∎
@@
Let $F$ be a decision forest on a set $V$ with $f$ as in Definition 1.3. Let $\mathcal{F}$ be the set of $(F,\supseteq)$’s connected components. Then, for any $v\in V$ and any $T\in\mathcal{F}$, we have $f(v)\subseteq T$ iff $v\in V_{T}$.
@@
Let $v\in V$ and $T\in\mathcal{F}$. By Proposition 1.4, $v\in V_{T}$ iff $V_{T}\in f(v)$. By Lemma A.47, Part 2, this is equivalent to $f(v)\subseteq T$. ∎
@@
(Ad necessity):  Suppose $F$ to be a decision forest on $V$. Let $W$ be the set of maximal chains in $(F,\supseteq)$, and $f\colon V\rightarrow W$ a bijection such that, for every $x\in F$, $({\mathcal{P}}f)(x)=W(x)$, just as in Definition 1.3.
We claim that for all $v\in V$ there is a unique $T\in\mathcal{F}$ such that $f(v)\subseteq T$. Property 1 then follows from Lemma A.48.
For the proof, let $v\in V$. As s maximal chain in the rooted forest $(F,\supseteq)$, $f(v)$ is non-empty. Hence, there is $x\in f(v)$. As $\mathcal{F}$ is a partition of $F$, there is a unique $T\in\mathcal{F}$ with $x\in T$. For arbitrary $y\in f(v)$, we have $x\subseteq y$ or $y\subseteq x$, since $f(v)$ is a chain. Hence, $x\sim y$, or in other words $y\in T$. Thus $T$ is the unique element of $\mathcal{F}$ with $f(v)\subseteq T$.
For Property 2, let $T\in\mathcal{F}$. Then $(T,\supseteq)$ defines a rooted tree with root $V_{T}$, by definition of $\mathcal{F}$. Moreover, for all $v\in V$, we have $f(v)\in W(V_{T})$ iff $v\in V_{T}$, by Proposition 1.4. Hence, $f_{T}:=f|_{V_{T}}$ yields a bijection $V_{T}\rightarrow W(V_{T})$.
By Lemma A.47, Part 3, second sentence, $W(V_{T})$ equals the set of maximal chains in $(T,\supseteq)$. By definition of $V_{T}$, all $x\in T$ satisfy $V_{T}\supseteq x$ so that the definition of $f$ yields:
@@
By Lemma A.47, Part 3, first sentence, $W(x)$ consists exactly of all maximal chains in $(T,\supseteq)$ containing $x$. Hence, $T$ is its own representation by decision paths.
(Ad sufficiency):  Let $F$ satisfy Properties 1 and 2 from the theorem. It remains to show that $F$ is its own representation by decision paths, i.e. Part 2 in Definition 1.3.
By Property 2 and Lemma A.47, Part 3, for each $T\in\mathcal{F}$, there is a bijection $f_{T}\colon V_{T}\rightarrow W(V_{T})$ such that for all $x\in T$, $(\mathcal{P}f_{T})(x)$ equals the set of decision paths in $(T,\supseteq)$ containing $x$. By Part 3 of Lemma A.47, this set is equal to $W(x)$. By Property 1, we obtain a map $f\colon V\rightarrow W$ satisfying, for any $T\in\mathcal{F}$ and $v\in V_{T}$, $f(v)=f_{T}(v)$. $f$ is injective by Lemma A.47, Part 2, since $\mathcal{F}$ is a partition. $f$ is also surjective, hence bijective, since $F\neq\emptyset$ and so every maximal chain $w\in W$ contains an element $x$ for which there is $T\in\mathcal{F}$ with $x\in T$. Hence, $w\in W(x)$ which is a subset of $W(V_{T})$ by Lemma A.47, Part 3. Hence, there is $v\in V_{T}$ with $f(v)=f_{T}(v)=w$.
Let $x\in F$. Then there is $T\in\mathcal{F}$ with $x\in T$. Then $V_{T}\supseteq x$. Hence, $({\mathcal{P}}f)(x)=({\mathcal{P}}f_{T})(x)$. By definition of $f_{T}$ and Part 3 of Lemma A.47, we have
@@
Lemma A.47, Part 3, implies that the latter set equals $W(x)$. Thus, $(\mathcal{P}f)(x)=W(x)$. ∎